OpenStudy (chris215):

To three decimal places, find the value of the first positive x-intercept for the function f(x) = 2cos(x + 3).

OpenStudy (chris215):

I got -1.429 but i think that's wrong

OpenStudy (anonymous):

it has to be wrong, since it is not positive

OpenStudy (anonymous):

\[2\cos(x+2)=0\\ \cos(x+2)=0\\ x+2=\frac{\pi}{2}\] i guess that gives you the negative answer right?

OpenStudy (anonymous):

try \[x+2=\frac{5\pi}{2}\]